OpenStudy (anonymous):

Using the disk method find the volume of the area bounded by y = the square root of (x - 1) y = (x-1)/2 rotated around the x axis. I'll show you what I wrote in the next answer.

OpenStudy (anonymous):

So what I did was first find the limits of integration by setting the two y values equal to each other. Like so. .5(x-1) - (x-1)^1/2 = 0 I got x = 5 and x =1 for my x values. So I set my upper limit to 5, and my lower limit to 1.

OpenStudy (anonymous):

I then wrote \[\int\limits_{1}^{5} \frac{ (x-1) }{ 2 }^{2} - (x-1)\] and took the anti derivative of that which got me \[(\frac{ x-1 }{ 6 })^{3} - \frac{ (x-1) }{ 2 }^{2}\] When I solved for that I got (.29-4) -(0) which gave me a negative volume. Which step did I do incorrectly?

OpenStudy (anonymous):

the integrals. take them separately and you'll see \[\int\limits_{1}^{5}[(x-1)^{1/2} - x + 1]dx\]

OpenStudy (anonymous):

Huh? Where did the .5? Go. The original equation was. \[\frac{ (x-1) }{ 2 }\] and \[\sqrt{x-1}\] I just rewrote them. Are you saying I should redo the integral step? Let's see what I get.

OpenStudy (anonymous):

my bad. add it

OpenStudy (anonymous):

I can just square both, it's part of the formula when using the disk method to find volume.

OpenStudy (anonymous):

\[∫\frac{ (x−1) }{ 2 } dx=\frac{ 1 }{ 2 } \left(\frac{ x^2 }{ 2 }−x \right)≠\frac{ 1 }{ 2 }(x^2−2x+1)\]

OpenStudy (anonymous):

sorry for earlier. forgot how to do the actual disk integration so i focused on that

OpenStudy (anonymous):

instead of this

OpenStudy (anonymous):

make that dne 1/4(x-1)^2

OpenStudy (anonymous):

I still have no clue what you are talking about or no clue where I made my mistake. It seems like everything you are saying has nothing to do at all with what I am asking.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

second integral

OpenStudy (anonymous):

\[\int\limits_{1}^{5}[\frac{ (x-1)^2 }{ 4 }−(x−1)]dx\]

OpenStudy (anonymous):

except not 1/2 lol my bad

OpenStudy (anonymous):

\[\int\limits_{}^{}(x-1)dx = \frac{ x^2 }{ 2 } - x \neq \frac{ 1 }{ 2 }(x^2 -2x + 1)\]

OpenStudy (anonymous):

So what you are trying to say is that the anti derivative of (x-1) is NOT (x-1)^2/2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well what is it then? Do you mean that I should not get rid of the square root untill after I have taken the anti derivative? If I don't get rid of the square root I get a non real answer.

OpenStudy (anonymous):

So iSo instead it would be 

OpenStudy (anonymous):

no just \[\int\limits_{}^{}(x-1)dx = \frac{ x^2 }{ 2 } - x \neq \frac{ (x-1)^2 }{ 2 }\] integral is fine